# Twin branching in shape memory alloys: a 1D continuum model with energy dissipation effects

### v1

@@ 3.4 Model predictions
Figure 4: Individual contributions to the total free energy $\Phi$ as a function of $\bar{\gamma}$: (a) dimensionless energy contributions $\bar{\Phi}_{i}$; (b) fractions $\Phi_{i}/\Phi$.
@@ 4.1 Overall dissipation potential
In the continuum setting, the twin spacing $h$ is now considered as a function of both position and time, $h=h(x,t)$. Additionally, the length of the twinned domain is time-dependent, such that $0\leq x\leq L(t)$.
@@
where $\hat{r}$ is the rate-independent threshold (critical driving force) for interface propagation, and the quantities with a hat refer to the local sharp-interface description. The dependence of the interface speed $\hat{v}_{\rm n}$ on the rate of change of the twin spacing, $\dot{h}$, is derived in A, see Eq. (42), and reads
@@ 4.2 Evolution problem
where the subdifferential $\partial_{\dot{h}}|\dot{h}|$ and the resulting inclusion are introduced in view of non-differentiability of the dissipation potential. Due to the non-smoothness, a direct computational treatment of the above minimization problem is not immediate, and the augmented Lagrangian method (Alart and Curnier, 1991) is used for that purpose. The details are provided in B.
The dimensionless form of the weak form (32) reads
$0\in\int_{0}^{1}\left(\bar{h}_{,\bar{x}}\delta\bar{h}_{,\bar{x}}-\frac{2\bar{\gamma}}{\bar{h}^{2}}\delta\bar{h}+\partial_{\stackrel{{\scriptstyle\mbox{\tiny$\circ$}}}{{h}}}|\!\stackrel{{\scriptstyle\mbox{\tiny$\circ$}}}{{h}}\!|\,\frac{\bar{r}}{\bar{h}}\delta\bar{h}\right){\rm{d}}\bar{x}+\bar{\Gamma}\,\delta\bar{h}\big{|}_{\bar{x}=1}\qquad\forall\,\delta\bar{h},$ (33)
where $\bar{r}/\bar{h}$ is the dimensionless threshold for interface propagation with
$\bar{r}=\frac{r}{A}=\frac{3(1-\lambda)\hat{r}}{2A},$ (34)
and $\delta\bar{h}$ is the test function replacing $\delta\dot{h}$ in the weak form (32). Note also that ${{h}}$ denotes the rate of the dimensionless twin spacing $\bar{h}$ evaluated at a constant $x$ (rather than at a constant $\bar{x}$). Considering that the domain $(0,L)$ evolves in time, the actual computations are carried out on a fixed domain, $0\leq\bar{x}\leq 1$, and on a fixed finite-element mesh. This simplifies enforcement of the boundary condition at $x=L$ ($\bar{x}=1$), but requires a proper computation of the rate of twin spacing. This is discussed further in B.
@@ 4.3 Model predictions
For the purpose of presentation of the results, it is convenient to introduce an alternative normalization based on the maximum length $L_{\rm max}$, rather than on the current length $L$ as in Eqs. 17, 19, and (20). The corresponding dimensionless quantities are denoted by a superimposed tilde, namely
@@
Sample results are presented in Fig. 5 for $L_{\rm max}=20\,\mu$m which corresponds to $\tilde{\gamma}=10^{-3}$ for the adopted material parameters. The evolution of twin spacing $\tilde{h}$ is shown in Fig. 5(a,b) where the individual curves correspond to selected values of the domain length $\tilde{L}$. As a reference, the dashed lines show the corresponding profiles of twin spacing obtained by minimization of the total free energy (no dissipation, $\hat{r}=0$). In the first stage of the evolution, when the domain length $L$ increases, Fig. 5(a), the twin spacing increases, however, the increase is slowed down as compared to the case with no dissipation. In the second stage, when the domain length $L$ decreases, Fig. 5(b), the evolution is more complex. Initially, there is a zone in which the twin spacing does not evolve. This, in particular, concerns the free end, at $x=0$, see Fig. 5(c). The evolution of the twin spacing at the other end of the domain, at $x=L$, is shown in Fig. 5(d).
@@
Figure 5: Evolution of twin spacing $\tilde{h}$ for $\tilde{\gamma}=10^{-3}$ ($L_{\rm max}=20\,\mu$m) and $\bar{r}=0.21$ ($\hat{r}=4$ MPa): (a,b) $\tilde{h}$ as a function of $\tilde{x}$ while $\tilde{L}$ is increasing (a) and decreasing (b); (c,d) twin spacing at $x=0$ (c) and at $x=L$ (d) as a function of the domain length $\tilde{L}$. The dashed lines correspond to the case with no dissipation, $\hat{r}=0$.
Fig. 6 shows the evolution of the total free energy $\tilde{\Phi}$ and its contributions defined in Eq. (23). While the effect of energy dissipation on the individual contributions is significant, see Fig. 6(b,c,d), the total energy is only marginally affected and is very close to the case of no dissipation (depicted by the dashed line), see Fig. 6(a). Clearly, in the dissipative case, the total free energy is higher than in the case with no dissipation, since the later case corresponds to the minimum of the total free energy.
@@
Figure 7: Evolution of the twin spacing $\tilde{h}$ at $x=0$ (a) and at $x=L$ (b), total free energy $\tilde{\Phi}$ (c) and its interfacial contribution $\tilde{\Phi}_{\rm int}$ (d) for $\tilde{\gamma}=10^{-2}$ ($L_{\rm max}=2\,\mu$m) and $\bar{r}=0.21$ ($\hat{r}=4$ MPa). The dashed lines correspond to the case with no dissipation, $\hat{r}=0$.
@@
Figure 8: Evolution of the twin spacing $\tilde{h}$ at $x=0$ (a) and at $x=L$ (b), total free energy $\tilde{\Phi}$ (c) and its interfacial contribution $\tilde{\Phi}_{\rm int}$ (d) for $\tilde{\gamma}=10^{-4}$ ($L_{\rm max}=200\,\mu$m) and $\bar{r}=0.21$ ($\hat{r}=4$ MPa). The dashed lines correspond to the case with no dissipation, $\hat{r}=0$.
@@ 5 Discussion
It is important to remark that there is quite some uncertainty concerning the values of the material parameters, particularly, the twin boundary energy $\gamma$ and the rate-independent interface propagation threshold $\hat{r}$. Also the value of parameter $A$ characterizing the elastic strain energy of branching has been determined using the upper bound estimate provided by Seiner et al. (2020). To allow direct comparison with the predictions of the model of Seiner et al. (2020), we have used the estimate corresponding to isotopic elastic properties, Eq. (4). On the other hand, it is well known that the CuAlNi alloy is elastically highly anisotropic (Sedlák et al., 2005), hence accounting for the elastic anisotropy may influence the value of $A$ (the corresponding estimate has also been provided by Seiner et al. (2020)). In any case, the actual accuracy of those estimates is not known. Alternative estimates, for instance, from direct simulations of branched microstructures, are not available.
@@ 6 Conclusion
A new approach to the modeling of twin branching in shape memory alloys has been developed in which twin spacing is treated as a continuous function of the distance from the austenite–martensite interface. The free energy, which comprises the interfacial and elastic strain contributions, is then expressed in terms of the twin spacing and its gradient. The total free energy is minimized, and the problem is solved numerically using the finite element method. The resulting 1D model can be considered a continuum counterpart of the discrete model of Seiner et al. (2020), and a very good agreement between the two models has been demonstrated within the entire range of physically relevant model parameters.
The dimensionless formulation of the model, which in the continuum setting is introduced in a natural way, indicates that the problem is fully characterized by two dimensionless parameters expressed in terms of the material properties and of the size of the twinned domain. The results show that, in a wide range of values of the dimensionless interfacial energy $\bar{\gamma}$, the solution is practically insensitive to parameter $\bar{\Gamma}$ characterizing the energy of elastic micro-strains at the A–MM interface. In this regime, the problem is thus fully characterized by only one dimensionless parameter, $\bar{\gamma}$.
The evolution of the branched microstructure has also been studied by considering the energy dissipation associated with propagation of the twin interfaces. The evolution is governed by a minimization problem formulated for a rate potential comprising the free energy rate and the dissipation potential. The results show that, for a high but realistic rate-independent threshold for interface propagation, the energy dissipation effects are significant only for high values of the dimensionless interfacial energy parameter $\bar{\gamma}$ (for instance, for a relatively small size of the twinned domain). Otherwise, the energy dissipation effects are negligible so that effectively the problem is history independent and minimization of the free energy alone suffices to determine the twin spacing and other characteristics. Interestingly, the twin spacing and the individual contributions to the total free energy exhibit a visibly higher sensitivity to the dissipation effects than the total free energy itself.
To the best of our knowledge, this is the first time that the energy dissipation effects are included in the modeling of twin branching phenomena. This has been facilitated by the continuum description adopted in this work, as otherwise modeling of an evolution problem would be difficult, if possible at all.
@@ Acknowledgments
This work has been partially supported by the National Science Centre (NCN) in Poland through the Grant No. 2021/43/D/ST8/02555. For the purpose of Open Access, the authors have applied a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
Appendix A Estimation of local interface propagation speed
In this appendix, an estimate of the local interface propagation speed $\hat{v}_{\rm n}$ is derived as a function of $\dot{h}$, the rate of change of the twin spacing. Referring to the discrete model of branching shown in Fig. 1, we assume that locally the evolution of the branched microstructure, i.e., the change of the twin spacing, proceeds by a ‘rigid’ translation of the cells, see Fig. 1. For the purpose of estimating the relationship between the propagation speed of individual interfaces and the rate of change of the twin spacing, we thus neglect the possible change of the proportions of the cells.

### v2

@@ 3.4 Model predictions
Figure 4: Individual contributions to the total free energy $\Phi$ as a function of $\bar{\gamma}$ (for $\bar{\Gamma}=0.35$): (a) dimensionless energy contributions $\bar{\Phi}_{i}$; (b) fractions $\Phi_{i}/\Phi$.
@@ 4.1 Overall dissipation potential
In the continuous setting, the twin spacing $h$ is now considered as a function of both position and time, $h=h(x,t)$. Additionally, the length of the twinned domain is time-dependent, such that $0\leq x\leq L(t)$.
@@
where $\hat{r}$ is the rate-independent threshold (critical driving force) for interface propagation, and the quantities with a hat refer to the local sharp-interface description. The dependence of the interface speed $\hat{v}_{\rm n}$ on the rate of change of the twin spacing, $\dot{h}$, is derived in B, see Eq. (52), and reads
@@
The above simple description of the dissipative effects is based on the assumption that the dissipation results from the propagation of twin interfaces only. However, the actual evolution of the branched microstructure may also be associated with abrupt evens, like formation or disappearance of individual twin lamellae. In agreement with the point of view adopted in the present continuous model, such events are assumed to appear locally at different locations and at different time instants, hence are not observed on the macroscale upon spatial and time averaging. Secondly, the energy dissipation associated with such abrupt events may require a different description than that corresponding to a smooth propagation of twin interfaces. In the present model, such effects can only be accounted for by adjusting the rate-independent threshold $\hat{r}$. A more refined description would be an interesting topic for future work.
@@ 4.2 Evolution problem
where the subdifferential $\partial_{\dot{h}}|\dot{h}|$ and the resulting inclusion are introduced in view of the non-differentiability of the dissipation potential. Due to the non-smoothness, a direct computational treatment of the above minimization problem is not immediate, and the augmented Lagrangian method (Alart and Curnier, 1991) is used for that purpose. The details are provided in C.
Let us now derive a dimensionless form of Eq. (34). As in Section 3.2, the twin spacing $h$ is normalized by the domain length $L$, thus $\bar{h}=h/L$, and the dimensionless twin spacing $\bar{h}=\bar{h}(\bar{x},t)$ is defined on a fixed domain, $0\leq\bar{x}\leq 1$. Since the evolution problem is rate-independent, the solution is invariant upon arbitrary scaling of the time, and hence the time $t$ is not normalized.
The time derivative of the twin spacing, $\dot{h}=\partial h/\partial t$, enters the weak form (34) through the subdifferential $\partial_{\dot{h}}|\dot{h}|$. To maintain the physical meaning of $\dot{h}$ in the dimensionless setting, the time derivative of the dimensionless twin spacing $\bar{h}$ must be taken at a constant $x$ rather than at a constant $\bar{x}$. The corresponding time derivative is denoted by ${\bar{h}}$ and satisfies
${\bar{h}}(\bar{x},t)=\frac{1}{L}\,\dot{h}(x,t)\big{|}_{x=L\bar{x}},$ (35)
see C for the computational treatment of this term.
The weak form (34) involves also the test function $\delta\dot{h}$ that has an interpretation of the variation of the twin spacing rate. It is an arbitrary function of sufficient regularity, just like $\delta h$ in Eq. (13). Accordingly, the test function $\delta\dot{h}$ can be equivalently replaced by $\delta h$ in the weak form (34), and subsequently $\delta\bar{h}=\delta h/L$ will be used as the test function in the dimensionless formulation of the weak form (34) which reads
$0\in\int_{0}^{1}\left(\bar{h}_{,\bar{x}}\delta\bar{h}_{,\bar{x}}-\frac{2\bar{\gamma}}{\bar{h}^{2}}\delta\bar{h}+\partial_{\overset{\mbox{\tiny$\circ$}}{\bar{h}}}\big{|}\overset{\mbox{\tiny$\circ$}}{\bar{h}}\big{|}\,\frac{\bar{r}}{\bar{h}}\delta\bar{h}\right){\rm{d}}\bar{x}+\bar{\Gamma}\,\delta\bar{h}\big{|}_{\bar{x}=1}\qquad\forall\,\delta\bar{h}.$ (36)
Compared to the static case in Eq. (13), the above weak form of the evolution problem involves an extra term introducing the dissipative driving force, where $\bar{r}/\bar{h}$ is the dimensionless threshold for interface propagation with
$\bar{r}=\frac{r}{A}=\frac{3(1-\lambda)\hat{r}}{2A}.$ (37)
The weak form (36) constitutes the basis for the finite-element treatment of the evolution problem. Despite the domain $0\leq x\leq L$ evolves in time, the actual computations are carried out on a fixed domain, $0\leq\bar{x}\leq 1$, and on a fixed finite-element mesh. This simplifies enforcement of the boundary condition at $x=L$ (at $\bar{x}=1$), but requires a proper computation of the rate of the twin spacing. This is discussed further in C.
@@ 4.3 Model predictions
For the purpose of presentation of the results, it is convenient to introduce an alternative normalization based on the maximum length $L_{\rm max}$, rather than on the current length $L$ as in Eqs. 18, 20, and (21). The corresponding dimensionless quantities are denoted by a superimposed tilde, namely
@@
Sample results are presented in Fig. 5 for $L_{\rm max}=20\,\mu$m which corresponds to $\tilde{\gamma}=10^{-3}$ for the adopted material parameters. The evolution of twin spacing $\tilde{h}$ is shown in Fig. 5(a,b) where the individual curves correspond to selected values of the domain length $\tilde{L}$. As a reference, the dashed lines show the corresponding profiles of twin spacing obtained by minimization of the total free energy (no dissipation, $\hat{r}=0$).
In the first part of the evolution, when the domain length $\tilde{L}$ increases (forward motion of the A–MM interface), the twin spacing increases, however, the increase is slowed down as compared to the case with no dissipation (dashed lines), see Fig. 5(a).
In the second part, when the domain length $\tilde{L}$ decreases (backward motion of the A–MM interface), Fig. 5(b), the evolution is more complex and comprises two stages. In the first stage of the backward motion, the twinned domain is divided into two zones. In the outer zone, closer to the A–MM interface at $\tilde{x}=\tilde{L}$, the twin spacing evolves, while in the inner zone, closer to the free end at $\tilde{x}=0$, the twin spacing does not evolve, as the evolution is stopped by the dissipative forces. This can be seen in Fig. 5(b) where the profiles of $\tilde{h}$ corresponding to $\tilde{L}=1$ and $\tilde{L}=0.8$ overlap over a part of the domain, approximately $0\leq\tilde{x}\leq 0.4$. The no-evolution zone gradually decreases during the first stage, and when it vanishes the second stage of the backward motion starts, in which the twin spacing evolves in the entire domain.
@@
Figure 5: Evolution of twin spacing $\tilde{h}$ for $\tilde{\gamma}=10^{-3}$ ($L_{\rm max}=20\,\mu$m) and $\bar{r}=0.21$ ($\hat{r}=4$ MPa): (a,b) $\tilde{h}$ as a function of $\tilde{x}$ while $\tilde{L}$ is increasing (a) and decreasing (b); (c,d) twin spacing at $\tilde{x}=0$ (c) and at $\tilde{x}=\tilde{L}$ (d) as a function of the domain length $\tilde{L}$. The dashed lines correspond to the case with no dissipation, $\hat{r}=0$.
The two-stage evolution of twin spacing during the backward motion can be seen in the plots of twin spacings $\tilde{h}_{0}$ and $\tilde{h}_{L}$, belonging to the ends, see Fig. 5(c,d). The transition from the first stage of the backward motion to the second one is marked by a kink in the evolution plots. In particular, at the free end, $\tilde{x}=0$, the twin spacing $\tilde{h}_{0}$ is constant during the first stage, i.e., until $\tilde{L}\approx 0.6$ is reached, see Fig. 5(c). When $\tilde{L}$ decreases further, during the second stage, the twin spacing $\tilde{h}_{0}$ starts to decrease. At the same time, the twin spacing at the A–MM interface at $\tilde{x}=\tilde{L}$, increases during the first stage of the backward motion and decreases during the second stage. Similar kinks are also seen in the plots of the evolution of the energy contributions that are discussed next.
Fig. 6 shows the evolution of the total free energy $\tilde{\Phi}$ and its contributions defined in Eq. (25). While the effect of energy dissipation on the individual contributions is significant, see Fig. 6(b,c,d), the total energy is only marginally affected and is very close to the case of no dissipation (depicted by the dashed line), see Fig. 6(a). Clearly, in the dissipative case, the total free energy is higher than in the case with no dissipation, since the later case corresponds to the minimum of the total free energy. As discussed above, the transition from the first stage of the backward motion of the A–MM interface to the second stage is associated with a kink at $\tilde{L}\approx 0.6$ in each of the curves in Fig. 6 (not visible in Fig. 6(a)).
@@
Figure 7: Evolution of the twin spacing $\tilde{h}$ at $\tilde{x}=0$ (a) and at $\tilde{x}=\tilde{L}$ (b), the total free energy $\tilde{\Phi}$ (c) and its interfacial contribution $\tilde{\Phi}_{\rm int}$ (d) for $\tilde{\gamma}=10^{-2}$ ($L_{\rm max}=2\,\mu$m) and $\bar{r}=0.21$ ($\hat{r}=4$ MPa). The dashed lines correspond to the case with no dissipation, $\hat{r}=0$.
@@
Figure 8: Evolution of the twin spacing $\tilde{h}$ at $\tilde{x}=0$ (a) and at $\tilde{x}=\tilde{L}$ (b), the total free energy $\tilde{\Phi}$ (c) and its interfacial contribution $\tilde{\Phi}_{\rm int}$ (d) for $\tilde{\gamma}=10^{-4}$ ($L_{\rm max}=200\,\mu$m) and $\bar{r}=0.21$ ($\hat{r}=4$ MPa). The dashed lines correspond to the case with no dissipation, $\hat{r}=0$.
@@ 5 Discussion
It is important to remark that there is quite some uncertainty concerning the values of the material parameters, particularly, the twin boundary energy $\gamma$ and the rate-independent interface propagation threshold $\hat{r}$. Also, the value of parameter $A$ characterizing the elastic strain energy of branching has been determined using the upper bound estimate provided by Seiner et al. (2020). To allow direct comparison with the predictions of the model of Seiner et al. (2020), we have used the estimate corresponding to isotropic elastic properties, Eq. (4). On the other hand, it is well known that the CuAlNi alloy is elastically highly anisotropic (Sedlák et al., 2005), hence accounting for the elastic anisotropy may influence the value of $A$ (the corresponding estimate has also been provided by Seiner et al. (2020)). In any case, the actual accuracy of those estimates is not known. Alternative estimates, for instance, from direct simulations of branched microstructures, are not available.
@@ 6 Conclusion
A new approach to the modeling of twin branching in shape memory alloys has been developed in which twin spacing is treated as a continuous function of the distance from the austenite–martensite interface. The free energy, which comprises the interfacial and elastic strain contributions, is then expressed in terms of the twin spacing and its gradient. The total free energy is minimized, and the problem is solved numerically using the finite element method. The resulting 1D model can be considered a continuous counterpart of the discrete model of Seiner et al. (2020), and a very good agreement between the two models has been demonstrated within the entire range of physically relevant model parameters.
The dimensionless formulation of the model, which in the continuous setting is introduced in a natural way, indicates that the problem is fully characterized by two dimensionless parameters expressed in terms of the material properties and of the size of the twinned domain. The results show that, in a wide range of values of the dimensionless interfacial energy $\bar{\gamma}$, the solution is practically insensitive to parameter $\bar{\Gamma}$ characterizing the energy of elastic micro-strains at the A–MM interface. In this regime, the problem is thus fully characterized by only one dimensionless parameter, $\bar{\gamma}$.
The evolution of the branched microstructure has also been studied by considering the energy dissipation associated with propagation of the twin interfaces. The evolution is governed by a minimization problem formulated for a rate potential comprising the free energy rate and the dissipation potential. The results show that, for a high but realistic rate-independent threshold for interface propagation, the energy dissipation effects are significant only for high values of the dimensionless interfacial energy parameter $\bar{\gamma}$ (for instance, for a relatively small size of the twinned domain). Otherwise, the energy dissipation effects are negligible so that effectively the problem is history-independent and minimization of the free energy alone suffices to determine the twin spacing and other characteristics. Interestingly, the twin spacing and the individual contributions to the total free energy exhibit a visibly higher sensitivity to the dissipation effects than the total free energy itself.
To the best of our knowledge, this is the first time that the energy dissipation effects are included in the modeling of twin branching phenomena. This has been facilitated by the continuous description adopted in this work, as otherwise modeling of an evolution problem may not be readily feasible.
@@ Acknowledgments
This work has been partially supported by the National Science Centre (NCN) in Poland through the Grant No. 2021/43/D/ST8/02555.
Appendix A Analytical solution for the static case
For completeness, an analytical solution for the static case presented in Section 3 is provided in this appendix. The derivation follows that reported by Zhang et al. (2024). The difference is that we exploit the boundary condition at $x=L$, Eq. (16)2, which is not considered by Zhang et al. (2024). Secondly, we provide a full final form without a simplifying assumption, as adopted in their work. This is commented at the end of the appendix.
The derivations are carried out in a dimensionless form. The starting point is the equation of equipartition of energy (17),
$\frac{1}{2}(\bar{h}_{,\bar{x}})^{2}-\frac{2\bar{\gamma}}{\bar{h}}=\bar{C}.$ (39)
We note that at $\bar{x}=0$ we have $\bar{h}_{,\bar{x}}=0$, cf. Eq. (19)1, which allows us to determine the constant $\bar{C}=-2\bar{\gamma}/\bar{h}_{0}$ so that Eq. (39) can be rewritten in the following form
$\frac{{\rm{d}}\bar{h}}{{\rm{d}}\bar{x}}=-2\sqrt{\bar{\gamma}\left(\frac{1}{\bar{h}}-\frac{1}{\bar{h}_{0}}\right)}.$ (40)
Here, we anticipate that the twin spacing is a monotonically decreasing function, hence the minus sign on the right-hand side.
Let us now consider $\bar{x}$ to be a function of $\bar{h}$ so that Eq. (40) implies
$\frac{{\rm{d}}\bar{x}}{{\rm{d}}\bar{h}}=-\frac{1}{2}\sqrt{\frac{\bar{h}_{0}\bar{h}}{\bar{\gamma}(\bar{h}_{0}-\bar{h})}}.$ (41)
This equation can be integrated analytically, thus yielding
$\bar{x}(\bar{h})=\frac{1}{2}\sqrt{\frac{\bar{h}_{0}^{3}}{\bar{\gamma}}}\left(\sqrt{\frac{\bar{h}}{\bar{h}_{0}}\left(1-\frac{\bar{h}}{\bar{h}_{0}}\right)}-\arcsin\sqrt{\frac{\bar{h}}{\bar{h}_{0}}}\right)+\bar{B}.$ (42)
The integration constant $\bar{B}$ can be determined from the condition that $\bar{x}(\bar{h}_{0})=0$, which gives
$\bar{B}=\frac{\pi}{4}\sqrt{\frac{\bar{h}_{0}^{3}}{\bar{\gamma}}}.$ (43)
Finally, we have
$\bar{x}(\bar{h})=\frac{1}{2}\sqrt{\frac{\bar{h}_{0}^{3}}{\bar{\gamma}}}\left(\sqrt{\frac{\bar{h}}{\bar{h}_{0}}\left(1-\frac{\bar{h}}{\bar{h}_{0}}\right)}+\frac{\pi}{2}-\arcsin\sqrt{\frac{\bar{h}}{\bar{h}_{0}}}\right).$ (44)
In this formula, $\bar{h}_{0}$ is unknown and can be determined from the boundary conditions at $\bar{h}=\bar{h}_{L}$, i.e., at the A–MM interface. Since $\bar{h}_{L}$ is also unknown, we have two unknowns ($\bar{h}_{0}$, $\bar{h}_{L}$) and two equations,
$\bar{x}(\bar{h}_{L})=1,\qquad\frac{{\rm{d}}\bar{x}}{{\rm{d}}\bar{h}}(\bar{h}_{L})=-\frac{1}{\bar{\Gamma}}.$ (45)
These equations must be solved numerically.
Note that Zhang et al. (2024) adopt a simplifying assumption concerning the integration constant $\bar{B}$, namely $\bar{B}\approx 1$, so that their Eq. (11) takes a particularly simple form. While this assumption may be a reasonable approximation in some cases, it does not hold in general, i.e., in the entire range of physically relevant values of the governing parameters $\bar{\gamma}$ and $\bar{\Gamma}$.
Appendix B Estimation of local interface propagation speed
In this appendix, an estimate of the local interface propagation speed $\hat{v}_{\rm n}$ is derived as a function of $\dot{h}$, the rate of change of the twin spacing. Referring to the discrete model of branching shown in Fig. 1, we assume that locally the evolution of the branched microstructure, i.e., the change of the twin spacing, proceeds by a ‘rigid’ translation of the cells, see Fig. 1. In this derivation, we neglect the possible change of the proportions of the cells.